OpenStudy (anonymous):

2x-y=12 write the equation in slope-intercept form then identify the slope and the y-intercept for each line.

OpenStudy (aroub):

-y=-2x+12 y=2x-12 slope=2 y-intercept=-12

OpenStudy (anonymous):

-y=12-2x, y=-12+2x y=2x-12